# Improving Retrieval-Augmented Generation through Multi-Agent Reinforcement Learning

### v1

@@ 4.4. Multi-Agent Optimization
for $epoch\leftarrow 1$ to $N\_epoch$ do
       for $batch\leftarrow 1$ to $N\_batch$ do
@@
                   Get sub-questions $subq$ for initial question $q$
@@
                   Retrieve $K$ candidate documents to construct $D$ // Selector (S)
@@
                   Select $IDs$ of helpful documents, and get $D_{\text{selected}}$
@@
The pseudocode for multi-agent optimization based on MAPPO is shown in Algorithm 1, which corresponds to the overall framework of MMOA-RAG depicted in Figure 1. For a specific question, the first step is to execute the Collect Rollout process. This process involves passing through the Query Rewriter, Retriever, Selector, and Generator, and the computed tuple $\mathcal{T}=\left((O_{QR},subq,R_{QR}),(O_{S},IDs,R_{S}),(O_{G},Ans_{\text{predict}},R_{G})\right)$ is stored in the replay buffer $\mathcal{M}$. Next, the Policy and Value Optimization process is executed where the GAE is used to estimate the advantage function $\hat{A}_{{\pi}_{\theta}}^{i,t}$. Subsequently, the overall loss function $\mathcal{L}(\theta,\phi)$ is calculated, and the parameters of both the Actor and Critic models are updated. Additionally, to accelerate the entire training process, we can run a minibatch in parallel. Ultimately, we obtain a well-trained Actor model used for subsequent inference and evaluation.
5. Experiments
  • RQ1. How does our MMOA-RAG perform compared to existing RAG optimization methods?
  • RQ2. What are the results of the ablation study on MMOA-RAG, specifically, is joint optimization of multiple modules more effective?
  • RQ3. Does the joint optimization method of MMOA-RAG exhibit generalizability across different RAG systems?
  • RQ4. How does the MMOA-RAG method perform in out-of-domain scenarios, i.e., what is its generalization capability?
5.1. Experimental Settings
5.1.1. Datasets and Evaluation
We conducted experiments using MMOA-RAG alongside various baseline models across three open-domain QA datasets: HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), and AmbigQA (Min et al., 2020). The candidate documents are all retrieved from Wikipedia passages for three datasets. We employ three key evaluation metrics—Accuracy, Exact Match (EM), and F1 score—to assess the performance of the RAG methods.
5.1.2. Baselines
@@
SELF-RAG (Asai et al., 2023): This innovative framework advances LLM performance by incorporating both adaptive retrieval mechanisms and self-reflection processes, aiming to produce precise and dependable answers.
RetRobust (Yoran et al., 2023): This approach fortifies the RAG architecture against irrelevant contexts, thereby boosting its effectiveness in open-domain question-answering scenarios.
Rewrite-Retrieve-Read (Ma et al., 2023): A small-scale query rewriter model is trained using reinforcement learning, optimizing the interaction between retrieval and answer generation.
BGM (Ke et al., 2024): Utilizing PPO, this method trains a bridge component to filter and identify documents that are more likely to be helpful, thus refining the quality of the retrieved context.
5.1.3. Implementation Details
We utilize Contriever (Izacard et al., 2021) as the Retriever. Regardless of how many sub-questions $subq$ the Query Rewriter generates from the initial question $q$, the Selector consistently receives a fixed set of $K=10$ documents as input. For example, if the Query Rewriter yields 2 sub-questions, each sub-question is used for retrieval, with the top-5 documents from each retrieval being selected as part of the candidate documents $D$ for the Selector. Furthermore, it is important to emphasize that we do not utilize any support facts or positive passages  [^3] Some QA datasets inherently include supportive texts that aid in answering questions. And some studies incorporate these supportive texts alongside retrieved candidate documents as input to LLMs for answer prediction, which significantly enhances answer quality. However, we adhere to the natural RAG process by using only the candidate documents provided by the retriever for answer prediction, as annotated supportive texts are not present in the natural RAG workflow. that come with the official datasets to generate answers. Instead, we only use the $K$ candidate documents from the retrieval model as external knowledge for answer generation.
Besides, we employ Llama-3-8B-Instruct (Dubey et al., 2024) as the foundational LLM for the baselines and MMOA-RAG. Building on the PPO code from LLama-Factory [^4] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024), we have developed MMOA-RAG, which optimizes the RAG multi-agent system using Multi-Agent PPO. And the critical hyperparameters of MMOA-RAG are detailed in Table 2.
Table 2. Key hyperparameters in the training process of MMOA-RAG.
@@ 5.1.3. Implementation Details
$lr$ Maximum learning rate 2e-5
@@
Table 3. Performance for different methods across datasets. All the results in this table are obtained using Llama-3-8B-Instruct as the backbone. In each dataset, the highest baseline value is underscored. The final row $\Delta$ displays the improvement of MMOA-RAG over the best baseline.
@@
LLM w/o RAG 25.08 21.31 31.18 27.78 23.68 29.47 27.21 20.96 33.42
Vanilla RAG w SFT 36.18 32.30 44.49 39.47 38.28 43.36 34.41 30.74 44.36
SELF-RAG (Asai et al., 2023) 30.42 27.77 38.93 36.32 35.39 38.86 28.35 25.70 39.04
RetRobust (Yoran et al., 2023) 37.69 34.60 46.49 41.02 39.73 44.51 35.13 32.37 44.78
Rewrite-Retrieve-Read (Ma et al., 2023) 38.03 33.93 46.32 40.40 39.17 44.17 35.94 31.90 45.92
BGM (Ke et al., 2024) 36.05 32.76 44.54 39.61 38.61 43.29 36.01 32.53 45.76
MMOA-RAG (ours) 39.15 36.15 48.29 42.73 41.52 46.40 38.85 34.75 48.59
$\Delta$ +1.12 +1.55 +1.80 +1.71 +1.79 +1.89 +2.84 +2.22 +2.67
5.2. Comparisons with Other Methods
We conducted a comparative analysis of MMOA-RAG against multiple baselines, with the results presented in Table 3. To ensure fairness in comparison, all baselines were re-implemented according to the settings delineated in Section 5.1.3. Each method utilized Llama-3-8B-Instruct as the backbone architecture. Within these methods, the untuned modules employed the pre-trained version of Llama-3-8B-Instruct, whereas the trainable modules were derived through specific SFT processes on Llama-3-8B-Instruct. Notably, in the Rewrite-Retrieve-Read framework, the query rewrite module, which is trainable, was optimized using the PPO algorithm applied to Llama-3-8B-Instruct; meanwhile, answer generation was implemented based on the SFT-refined backbone. Regarding the BGM method, we rebuilt the bridge to connect the retrieval model and the generation model, leveraging Llama-3-8B-Instruct, with this bridge being trained using the PPO algorithm. The generation model for BGM was similarly obtained from the SFT-refined backbone.
The results in Table 3 and the analysis in Section 5.2 jointly answer the RQ1.
5.3. Ablation Experiments on the Optimization of Different Agents
[…]
Methods HotpotQA 2WikiMultihopQA AmbigQA
Acc EM F1 Acc EM F1 Acc EM F1
MMOA-RAG w/o G 38.28 35.19 46.91 41.93 40.80 45.62 37.92 33.65 47.77
MMOA-RAG w/o S 38.81 35.27 47.61 42.47 40.99 45.95 38.69 34.38 48.23
MMOA-RAG w/o QR 37.26 34.66 46.05 42.20 40.82 45.70 38.64 34.34 48.14
MMOA-RAG 39.15 36.15 48.29 42.73 41.52 46.40 38.85 34.75 48.59
Table 5. Generality Experiments on RAG Systems with Varying Module Configurations. In the second column, SFT and MAPPO refer to the current module configuration following the warm start training stage (Section 4.3) and the MAPPO joint optimization training stage (Section 4.4), respectively. The symbol $\Delta$ signifies the enhancement achieved in the MAPPO stage relative to the SFT stage.
Modules Training Stage & Delta HotpotQA 2WikiMultihopQA AmbigQA
Acc EM F1 Acc EM F1 Acc EM F1
QR+S+G SFT 36.00 33.04 44.69 39.54 38.50 42.97 36.55 32.60 46.71
MAPPO 39.15 36.15 48.29 42.73 41.52 46.40 38.85 34.75 48.59
$\Delta$ +3.15 +3.11 +3.60 +3.19 +3.02 +3.43 +2.30 +2.15 +1.88
S+G SFT 34.25 32.18 43.14 38.93 37.97 42.40 35.85 32.35 45.82
MAPPO 38.23 34.85 47.07 41.79 40.57 45.25 37.60 33.90 47.19
$\Delta$ +3.98 +2.67 +3.93 +2.86 +2.60 +2.85 +1.75 +1.55 +1.37
QR+G SFT 36.76 32.78 45.00 39.15 37.89 42.91 35.50 31.50 45.31
MAPPO 38.90 35.89 47.94 42.43 41.01 46.19 37.65 33.50 47.53
$\Delta$ +2.14 +3.11 +2.94 +3.28 +3.12 +3.28 +2.15 +2.00 +2.22
As illustrated in Table 4, the complete version of MMOA-RAG, where all three modules are jointly optimized, delivers the highest performance. This underscores the effectiveness of multi-agent joint optimization within the RAG system and validates the importance of optimizing multiple modules concurrently. Additionally, the MMOA-RAG w/o S variant achieves the best performance among the three ablation configurations. The Selector’s primary function is to refine the candidate document set $D$, yielding a higher-quality subset $D_{\text{selected}}$, which enhances the Generator’s ability to produce a superior answer $Ans_{\text{predict}}$. However, through the joint optimization process with MAPPO, the Generator acquires some denoising capabilities. Consequently, satisfactory results can be achieved even when the Selector is not optimized during joint optimization.
Figure 3. Ablation experiments on AmbigQA dataset. The horizontal axis represents the number of training samples, while the vertical axis denotes the shared reward $R_{\text{shared}}$ (F1 score) during the training process.
We also present the trajectory of the shared reward $R_{\text{shared}}$ during the training process based on ablation experiments conducted on the AmbigQA dataset, as shown in Figure 3. From Figure 3, it is evident that the reward curve for MMOA-RAG demonstrates the fastest convergence rate and achieves the highest final convergence value. This underscores the effectiveness of joint optimization across multiple modules in significantly and efficiently enhancing the performance of the RAG system. Furthermore, the training curve for MMOA-RAG w/o G in Figure 3 is noticeably slower compared to other algorithms, and the test results for MMOA-RAG w/o G on the AmbigQA dataset, as shown in Table 4, are the poorest. These findings suggest that the Generator module is the most critical component for the single-hop AmbigQA dataset.
The results in Table 4 and Figure 3 answer the RQ.2 that it is more effective to optimize multiple modules in a RAG system simultaneously.
5.4. Generality Experiments on RAG Systems with Varying Module Configurations
5.5. Out-of-Domain Experiments
@@
Table 6. Out-of-domain experimental results: The model is trained on the HotpotQA dataset and tested on the AmbigQA dataset.
@@
SELF-RAG (Asai et al., 2023) 26.70 24.25 36.38
RetRobust (Yoran et al., 2023) 34.19 31.75 44.08
Rewrite-Retrieve-Read (Ma et al., 2023) 33.91 30.73 43.61
BGM (Ke et al., 2024) 32.58 29.77 42.07

### v2

@@ Appendix C The Pseudocode of Multi-Agent Training Process of MMOA-RAG
for $epoch\leftarrow 1$ to N_epochN\_epoch do
 for $batch\leftarrow 1$ to N_batchN\_batch do
@@
         Get sub-questions subqsubq for initial question $q$
@@
         Retrieve $K$ candidate documents to construct $D$
       // Selector (S)
@@
         Select IDsIDs of helpful documents, and get $D_{\text{selected}}$
@@
Appendix D The Detailed Introduction of the Implementation Details
D.1 Baselines
@@
SELF-RAG [1]: This innovative framework advances LLM performance by incorporating both adaptive retrieval mechanisms and self-reflection processes, aiming to produce precise and dependable answers.
RetRobust [52]: This approach fortifies the RAG architecture against irrelevant contexts, thereby boosting its effectiveness in open-domain question-answering scenarios.
Rewrite-Retrieve-Read [29]: A small-scale query rewriter model is trained using reinforcement learning, optimizing the interaction between retrieval and answer generation.
BGM [22]: Utilizing PPO, this method trains a bridge component to filter and identify documents that are more likely to be helpful, thus refining the quality of the retrieved context.
RAG-DDR [27]: This approach utilize DPO to align data preferences between different modules, enhancing performance and reducing hallucinations in RAG.
D.2 Implementation Details
To ensure fairness in comparison, all baselines were re-implemented using Llama-3-8B-Instruct as the backbone architecture. Within these methods, the untuned modules employed the pre-trained version of Llama-3-8B-Instruct, whereas the trainable modules were derived through specific SFT processes on Llama-3-8B-Instruct. Notably, in the Rewrite-Retrieve-Read framework, the query rewrite module, which is trainable, was optimized using the PPO algorithm applied to Llama-3-8B-Instruct; meanwhile, answer generation was implemented based on the SFT-refined backbone. Regarding the BGM method, we rebuilt the bridge to connect the retrieval model and the generation model, leveraging Llama-3-8B-Instruct, with this bridge being trained using the PPO algorithm. The generation model for BGM was similarly obtained from the SFT-refined backbone.
We utilize Contriever [16] as the Retriever. Regardless of how many sub-questions subqsubq the Query Rewriter generates from the initial question $q$, the Selector consistently receives a fixed set of $K=10$ documents as input. For example, if the Query Rewriter yields 2 sub-questions, each sub-question is used for retrieval, with the top-5 documents from each retrieval being selected as part of the candidate documents $D$ for the Selector. Furthermore, it is important to emphasize that we do not utilize any support facts or positive passages  [^3] Some QA datasets inherently include supportive texts that aid in answering questions. And some studies incorporate these supportive texts alongside retrieved candidate documents as input to LLMs for answer prediction, which significantly enhances answer quality. However, we adhere to the natural RAG process by using only the candidate documents provided by the retriever for answer prediction, as annotated supportive texts are not present in the natural RAG workflow. that come with the official datasets to generate answers. Instead, we only use the $K$ candidate documents from the retrieval model as external knowledge for answer generation.
Besides, we also employ Llama-3-8B-Instruct [7] as the foundational LLM for the baselines and MMOA-RAG. Building on the PPO code from LLama-Factory [^4] https://github.com/hiyouga/LLaMA-Factory [58], we have developed MMOA-RAG, which optimizes the RAG multi-agent system using Multi-Agent PPO. And the critical hyperparameters of MMOA-RAG are detailed in Table 6.
Table 6: Key hyperparameters in the training process of MMOA-RAG.
@@ D.2 Implementation Details
lrlr Maximum learning rate 2e-5
@@
Appendix E The Performance of Different Methods Based on Retriever BGE and E5
We firstly perform the training of different methods based on retrieval model Contriever [16], and the results are shown in Table 1.
We further test the trained models on different retriever BGE [46] and E5 [48]. From Table 7, we can see that our MMOA-RAG also surpasses most of baselines on each datasets.
Table 7: The performance of different methods based on Retriever BGE [46] and E5 [48]. Llama-3-8B-Instruct and its the fine-tuned version is the backbone for all methods. The highest baseline value in each dataset is underscored. The $\Delta$ displays the improvement of MMOA-RAG over the best baseline.
@@ Appendix E The Performance of Different Methods Based on Retriever BGE and E5
Retriever: BGE [46]
Vanilla RAG w/o train 37.72 28.99 41.62 28.86 18.54 27.87 48.05 35.10 52.31
Vanilla RAG w SFT 45.47 41.16 54.32 42.79 41.50 46.70 47.10 42.45 57.41
Self-RAG [1] 33.41 30.59 42.12 37.39 36.41 40.91 30.00 27.00 40.77
RetRobust [52] 46.66 42.74 55.77 47.03 45.57 50.62 47.00 42.20 58.14
Rewrite-Retrieve-Read [29] 45.58 40.95 54.35 43.91 42.48 47.83 46.30 41.75 57.19
BGM [22] 45.50 41.49 54.73 41.88 40.67 45.53 48.15 42.70 58.61
RAG-DDR [27] 45.13 41.51 54.32 42.49 41.21 46.00 48.50 42.95 58.71
MMOA-RAG 47.22 43.46 56.45 47.14 45.47 50.94 48.65 43.45 59.10
$\Delta$ +0.56 +0.72 +0.68 +0.11 -0.10 +0.32 +0.15 +0.50 +0.39
Retriever: E5 [48]
Vanilla RAG w/o train 36.92 28.01 40.80 28.57 17.50 27.67 50.50 37.00 53.93
Vanilla RAG w SFT 46.48 42.08 55.51 44.74 43.32 48.66 48.20 43.40 58.53
Self-RAG [1] 33.64 30.84 42.36 38.39 37.38 42.03 30.10 27.20 40.83
RetRobust [52] 46.66 42.78 55.85 49.27 47.59 52.75 48.85 44.10 59.29
Rewrite-Retrieve-Read [29] 46.62 42.00 55.54 46.41 44.83 50.35 47.85 43.55 58.36
BGM [22] 45.81 41.72 55.32 43.24 41.99 46.83 48.35 44.05 59.04
RAG-DDR [27] 44.50 41.09 53.96 43.89 42.49 47.29 49.60 43.75 59.75
MMOA-RAG 47.46 43.73 57.23 49.32 47.53 53.13 50.45 44.80 60.80
$\Delta$ +0.80 +0.95 +1.38 +0.05 -0.06 +0.38 -0.05 +0.70 +1.05
Appendix F Out-of-Domain Experiments
@@
Table 8: Out-of-domain experimental results: The model is trained on the HotpotQA dataset and tested on the AmbigQA dataset.
@@
SELF-RAG [1] 26.70 24.25 36.38
RetRobust [52] 34.19 31.75 44.08
Rewrite-Retrieve-Read [29] 33.91 30.73 43.61
BGM [22] 32.58 29.77 42.07